Gerry is to memorize 150 terms for a quiz. If the rate by which the number of terms Gerry memorizes is directly proportional to the number of terms already memorized and the number of remaining terms; and that initially Gerry knows 5 terms and 20 terms 5 minutes later. At what time is he half done?


1
Expert's answer
2021-12-15T10:19:37-0500

"\displaystyle
\text{Let m be the number terms Gerry memorizes, therefore}\\
\frac{dm}{dt}=km\\
\implies \frac{dm}{m}=kdt\\
\text{Integrating, we have that}\\
\ln m = kt + c\\
\text{Taking the exponential of both sides}\\
m = Ae^{kt}\\
\text{At t = 0, m =5}\\
A= 5\\
\implies m = 5 e^{kt}\\
\text{At t = 5, m =20}\\
m = 5e^{5k}\\
e^{5k}=4\\
\implies k = \frac{\ln 4}{5} = 0.277\\
75 = 5e^{0.277t}\\
\implies t = 9.776 \approx 10 min"